If 2.0 C of charge passes through a wire in time 0.25 s, what is the current in the wire? 8.0 V O 0.50 A O 0.50 V O 8.0 A Suppose that a current of 1.5 A is used to charge a conductor. After 4.0 s, how much charge is delivered to the conductor. 6.0 V O 6.0 C 2.7 C O 2.7 V
Answer #1

here,

1)

charge , Q = 2 C

time interval , t = 0.25 s

the current in the wire , I = Q /t

I = 2/0.25 A = 8 A

2)

current , I = 1.5 A

t = 4 s

the charge stored , Q = I * t

Q = 1.5 * 4 C = 6 C